OpenStudy (calculusxy):

Scientists are experimenting with a king of gun that may eventually be used to fire payloads directly into orbit. In one test, this gun accelerates a 5.0kg projectile from rest to speed of 4.0 x 10^3 m / s. The net force accelerating the projectile is 4.9 x 10^5N. How much time is required for the projectile to come up to speed?

OpenStudy (calculusxy):

@hartnn

hartnn (hartnn):

we have Force and Mass, so we can get acceleration The famous formula F= Ma

hartnn (hartnn):

then we can use the kinematic equation v = u+ at to get the time

OpenStudy (calculusxy):

So we'll only be using 4.9 x 10^5 and 5.0 kg?

hartnn (hartnn):

yep 4.9 * 10^5 = 5*a

hartnn (hartnn):

we'll be using this: "rest to speed of 4.0 x 10^3 m / s" to get the time initial velocity = u = rest = 0 speed

OpenStudy (anonymous):

This is deep man! :P, It's been long since I did this kind of math :)

hartnn (hartnn):

me too Sandeep :P I am googling the formulae ':)

OpenStudy (calculusxy):

Okay so for 4.9 * 10^5 = 5*a we have: 490,000 = 5a a= 98000

OpenStudy (anonymous):

@hartnn hehe +1

hartnn (hartnn):

yes, and now use the formula v = u+at "rest to speed of 4.0 x 10^3 m / s" initial velocity = u = rest = 0 speed final velocity = v = 4*10^3 a = 98000 find t

OpenStudy (calculusxy):

\[4000 = 0 + 98000t\] \[4000 = 98000t\]

OpenStudy (calculusxy):

t = 0.041

hartnn (hartnn):

seems correct :)

OpenStudy (calculusxy):

So that's my answer?

hartnn (hartnn):

in seconds, yes

OpenStudy (calculusxy):

Thanks!! Now last question.